### v8

# Increasing Trust in Language Models through the Reuse of Verified Circuits

Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of LMs built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of LMs.
Increasing Trust in Language Models through the Reuse of Verified Circuits
Philip Quirkethanks:    Core research contributor., Clement Neo†‡, Fazl Barezfootnotemark: †‡ Apart Research, University of Oxford
1 Introduction
Figure 1: An overview of our methodology: (1) We trained an accurate 6-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior. Hendrycks and Mazeika (2022) argue that making these models interpretable is key to their safe use. Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022). This understanding aids in predicting model behavior in new situations and fixing model errors.
In creating and training a model, we aim for high accuracy and trustworthiness. We achieve this by holding the model to a standard we term known-good. We define a model performing a task to be known-good if:
  1. 1.
    The model’s algorithm for the task and the mechanisms it implements (the “circuits") are understood.
  2. 2.
    All possible task edge cases have been identified and tested.
  3. 3.
    Empirically, the model prediction accuracy is 99.9999% (a standard reliability measure used in industry and abbreviated as “six nines"). That is, it can perform the task one million times with at most 1 wrong prediction.
Exhaustive testing of a model task may be infeasible. For instance, when adding two 5-digit integers (e.g. 12345+67890) there are ten billion variations. Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage all known edge cases. A known-good model must have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence. circuits that perform the task accurately.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and has six nines (99.9999%) accuracy. The model hence achieves our known-good standard. Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations, facilitating the interpretation of the model’s algorithm. We make progress toward a known-good model for both addition and subtraction.
Hence, our main contributions are three-fold:
  • Defining several known-good n-digit addition models with six nines accuracy which all use the same algorithm.
  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the new model’s algorithm.
  • Defining several n-digit addition and subtraction models with six nines accuracy, that reuse established addition model circuits for both operations, and detailing progress towards these models being known-good.
2 Related Work
Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020). Mathematical frameworks (Elhage et al., 2021a) explain how transformer attention heads can work with each other to implement complex algorithms.
Causal Scrubbing (Jenner et al., 2023) recommends explaining a model algorithm by documenting a low-level computation graph, mapping from the graph to the model nodes that implement the computation, and performing experimentation verification. Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems. It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu and Low, 2023).
Studying Toy Models for Arithmetic. Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023). Quirke and Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on rare edge cases (e.g. “77778+22222=100000" where a “carry 1" cascades through 4 digits), highlighting the importance of understanding and testing all edge cases for trustworthiness.
Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized" into discrete chunks (quanta) (Michaud et al., 2023). Models must learn these quanta to reduce loss. Understanding a network reduces to enumerating its quanta. Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
3 Methodology
Transformer models may learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our model implements this approach.
3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. We define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. Figure 2 shows an illustrative example.
Figure 2: For 5-digit addition, our model has 12 input (question) and 6 output (answer) token positions. We name the question tokens D4, …, D0, and D’4, …, D’0 and the answer tokens A5, …, A0. For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
First, we adopt the framework from Quirke and Barez (2024) for our model’s addition process. The “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of digit pairs, is defined as:
$A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The “Simple Carry" sub-task $A_{n}.SC$ determines whether the addition creates a carry bit:
$A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
While Quirke et al.’s model is capable of handling simple addition and carry bits generated directly from digit pair addition, it encounters difficulties with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
Consider “00144+00056=000210". Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position. A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110". The Quirke et al. 1-layer model could cascade carry bits across two digits, but not three or more.
3.2 Extending the Mathematical Framework
To answer “44444+55556=" with “100000", an accurate model must predict the first answer digit A5 as “1". To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits. This is especially hard as the model predicts answer tokens from left to right.
We introduce a digit-level sub-task called TriCase that calculates the essential “carry one" information for a single pair of digits $D_{n}$ and $D^{\prime}_{n}$. TriCase has 3 possible outputs representing a definite carry one (ST10), a possible carry one depending on the results of other calculations (ST9), or definitely not a carry one (ST8):
$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8,\\ \end{cases}\end{split}$ (3)
To perform the “cascading carry one" calculation, we introduce a TriAdd sub-task. It handles the case where a possible carry one becomes a definite carry one because the next lower digit pair generated a carry one. TriAdd is defined as:
$\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
The model can use $ST$ and $SV$ to accurately calculate A5 as 1 or 0 by using A5 = A4.SV = TriAdd(A4.ST, TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST)))). Note that in calculating A5, the model has also calculated an accurate carry bit for each answer digit. For example, the carry bit for A4 is A3.SV = TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST))).
With this framework, the model only needs the sub-tasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform addition (but some models also use the redundant $A_{n}$.SC sub-task). This framework, if implemented by a model, is sufficient for the model to perform n-digit addition accurately.
Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke et al. algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
Figure 3: To predict with 99.9999% accuracy, the addition algorithm first calculates “carry one" values $(A_{n}.ST)$, combining them into “cascading carry one" values $(A_{n}.SV).$ At the “+" token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0. The other answer digits are calculated by combining “base add" $(A_{n}.BA)$ and “carry one" $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
3.3 Techniques
In investigating the model circuits, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework. Hence, we define a node as the computation done by an attention head or MLP layer for a given token position. To investigate what each node is doing, we use the following techniques:
  1. 1.
    Intervention Ablation. To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across a batch and measure how that impacts downstream performance. We also use (automated, n-digit) intervention ablation tests, specific to each sub-task, to test for the expected sub-task behavior.
  2. 2.
    Attention Patterns. To find out what the model attends to at a node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
  3. 3.
    Principal Component Analysis (PCA). We use PCA to investigate the outputs of attention heads, especially where our framework suggests the head output may be tri-state or bi-state.
  4. 4.
    Question Complexity. We categorized questions by computational complexity (App. G). Addition categories (S0-S4) and subtraction categories (M0-M4) reflect the number of sequential digits a “carry one" or “borrow one" cascades through, respectively. We analyzed which nodes were necessary for correct predictions in each category.
4 Experiments
4.1 Training a Five-Digit Addition Model
The Quirke et al. 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%. Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App. D for alternatives tested). This configuration effectively doubled the computational power compared to the 1-layer model (see App. B for mode model configuration details). Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
We trained a 5-digit, 2-layer, 3-head model, with a 14 token vocabulary ( 0, .., 9, +, -, =, *, / ), batch size of 64, learning rate of 0.00008 and weight decay of 0.1. Training used an infinite dataset enriched with rare edge cases. Loss was defined as the mean across all answer tokens of their negative log likelihood loss. After 30 thousand training batches, the model’s final training loss was $\sim 2.3\times 10^{-8}$. Testing showed this model has six nines accuracy. (More details in App. C and Tab. 6).
(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
D’4 D’1 D’0 = + A5 A4 A3 A2 A1
L0H0 A4 A2 A1
L0H1 A5 A5..3 A5 A3 A0
L0H2 A5 A5..1
L0MLP A5..2 A4 A3 A2 A1 A0
L1H0
L1H2 A4 A3 A2
L1MLP A1 A0
Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits. After the "=" token, all nodes in a given token position are used to predict a single answer digit.
4.2 Investigating Five-Digit Addition
Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7). Further ablation experiments show that for these nine token positions, the model uses 36 nodes in predictions. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node. For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
The algorithm predicts the first answer digit, A5, at position P11. A5, which is always 0 or 1, is the most challenging to predict as it may rely on a long carry one cascade (e.g. 55555+44445=100000). An accurate algorithm must compute this cascade using the nodes located in positions P8 to P11. As illustrated in Figure 7, these nodes attend to all digit pairs from D4 D’4 to D0 D’0. Additionally, the PCA data, as shown in Figures 8 and 9, suggest that these nodes produce tri-state outputs. After our first two algorithm hypothesises failed testing (see App. H and I), we discovered that the model utilizes a minimal set of “carry one” information, leading to the development of the TriCase quanta. The model performs $A_{n}.ST$ using bigrams (see App. F) to map two input tokens to one result token e.g. “6” + “7” = ST10. In positions P8 to P11, the model does $A_{n}.ST$ calculations on all digit pairs from D4 D’4 to D0 D’0.
An MLP layer can be thought of as a “key-value pair" memory (Meng et al., 2022; Geva et al., 2021) that can hold many bigrams and trigrams. We posit our MLP implements the TriAdd function using bigrams and trigrams to calculate $A_{n}.SV$ values from $A_{n}.ST$ values.
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention. Figure 3 diagrams the algorithm, with node details in App. J. The model adheres to all known constraints and achieves six nines accuracy. We concluded this model instance is well-understood, well-functioning, and hence known-good.
The 1-layer model uses 21 nodes to achieve two nines (99%) accuracy. This model uses 36 nodes (an increase of 71%) to achieve six nines accuracy.
(P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’1 D’0 = + A6 A5 A4 A3 A2 A1
L0H0 A2.ST A3.ST A1.ST A4.ST A4.SC A3.SC A2.SC A1.SC A0.SC
L0H1 A1.ST A0.ST A5.SA A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST
Table 2: All addition models studied implement our addition algorithm. The algorithm SA, SC and $ST$ sub-tasks all exist for each digit and in appropriate token positions. For a sample model, this map shows the subtask locations. Interestingly, here each SA sub-task is shared across two attention heads.
4.3 Training n-digit Addition models
To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g. 111111+222222=+0333333).
These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
4.4 Investigating n-digit Addition models
We developed a declarative method to outline each sub-task. For instance, a sub-task may focus on question digits D2 and D’2, affect answer digit A3, influence S0 but not S1 complexity questions, and have specific PCA results and ablation tests. Using this declaration, we identified nodes performing these sub-tasks across all models.
An algorithm hypothesis, such as the one in section 4.2, is described by the required sub-tasks and their relationships. For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ sub-tasks for each question digit before the “=” token. We created a framework for declaring n-digit algorithm hypotheses and testing them against the sub-tasks found in each model, mapping the results (see Table 2 for an example).
Our seven 2-layer addition models all implement our addition algorithm. Given their six nines accuracy and implementation of the same algorithm, we can confirm these models as known-good.
4.5 Training n-digit Mixed models
To explore reuse, we initialized untrained models with a known-good addition-only model, then trained them to perform both subtraction and addition. We call these “mixed" models.
Specifically, we trained seven larger (6- or 10-digit, 2- or 3-layer, 3- or 4-head) models after initializing them with the weights from a known-good 2-layer 3-head addition model. The first 2 layers, first 3 heads of the mixed model were initialized with the addition model weights. We trained the mixed model with 80% subtraction and 20% addition batches. We enriched the (infinite) training dataset with rare addition and subtraction edge cases.
Some models achieved six nines accuracy and the others five nines. Attempts to “freeze" the inserted attention heads and/or MLP layers by periodically copying the addition weights back into the mixed model every 100 training steps resulted in lower accuracy. (Refer App.M and Tab.6)
Figure 4: Each mixed model was initialized with the weights from a known-good addition model, then was trained on 80% subtraction and 20% addition batches. A sample log loss graph (final loss 8.0e-9) is shown.
4.6 Investigating n-digit Mixed models
Unlike addition, subtraction question answers can be either positive or negative. Similar looking positive-answer (e.g. 10009-10000=+000009) and negative-answer (e.g. 10009-20000=-009991) subtraction questions can give answers that differ at several digit positions. We posited that the model treats three distinct question classes “addition", “positive-answer subtraction" and “negative-answer subtraction" differently. Ablation of nodes showed that some nodes only help predict one question class, some help predict two classes and some all three classes. It also showed that the inserted addition nodes are heavily used and the majority become polysemantic, performing both addition and subtraction calculations. (Details in Tab.3.)
Used Inserted
Question class # % # %
All questions 96 48
Addition 61 64% 42 88%
Positive-answer sub 70 73% 40 83%
Negative-answer sub 53 55% 29 60%
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction sub-tasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
To investigate the subtraction algorithm, and paralleling addition algorithm subtasks, we defined subtraction-specific sub-tasks Base Diff and Borrow One (see Tab. 4) and complexity measures (see App. G and  N).
Name +ve Sub -ve Sub Definition
Base Diff MD ND $D_{n}$- $D^{\prime}_{n}$10
Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
Table 4: We define 2 “positive-answer subtraction" and 2 “negative-answer subtraction" sub-tasks that parallel the addition sub-tasks Base Add SA and Carry One SC
We found that the inserted nodes that performed SA in the addition model, perform SA, MD and ND in the mixed model. (Refer Table 5.) These three sub-tasks are similar in that each performs a mapping from 100 input cases (10 $D_{n}$x 10 $D^{\prime}_{n}$) to 10 output cases (0..9). The mixed model “upgraded" these nodes to be polysemantic during training. Similarly, some SC addition nodes became polysemantic and now process SC, MB and/or NB .
(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
L0H0 A4.MT A3.MT A4.ST A2.SC A1.SC A0.SC OPR
A4.GT A3.MT A4.MT A4.SC A3.SC A2.NB A1.MB A0.MB SGN
OPR A1.NB A0.NB
L0H1 A2.ST A1.ST A3.ST A0.ST A5.SA
A4.ST A2.MT A1.MT A3.GT A0.MT A5.MD
A2.GT A1.GT A0.GT A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST A5.SA A4.MD A3.MD A2.MD A1.MD A0.MD
OPR A5.MD A4.ND A3.ND A2.ND A1.ND A0.ND
SGN A5.ND
L0H3 OPR OPR OPR OPR OPR OPR
SGN SGN SGN SGN SGN SGN
Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N. Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token. The MLP layers then select the output appropriate for the class. In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
An accurate subtraction model must answer “cascading borrow one" questions like 100000-000001=+0099999. We define the “essence of borrow one" subtask $MT$ used by both positive-answer (M) and negative answer (N) subtraction questions. $MT$ is like addition’s $ST$ subtask:
$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one. For N questions, the interpretation is the opposite.
Paralleling addition’s “cascading carry one" $A_{n}.SV$ calculation, we define “cascading borrow one" calculation sub-tasks $A_{n}.MV$ for positive-answer subtraction and $A_{n}.NV$ for negative-answer subtraction. We posit the MLP implements TriAdd-like functions using bigrams and trigrams to calculate $A_{n}.MV$ and $A_{n}.NV$ values from $A_{n}.MT$ values.
4.7 Mixed model question class detection
We posit that there is a specific circuit to detect whether a question is in the S, M or N class. If the question operator is “+" then the class is S, but if the question operator is “-" then the model must calculate if $D>=D^{\prime}$ to distinguish between the M and N classes. For accuracy, it needs this answer by the “=" token to predict the first answer token (the answer sign) as “+" or “-".
$D>=D^{\prime}$ can be derived from the $A_{n}.MT$ data. Alternatively, this calculation could be a distinct circuit using a new sub-task we define as:
$A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
For a 4 digit question, $D>=D^{\prime}$ can be calculated as A3.GT=1 or (A3.GT=0 and (A2.GT=1 or (A2.GT=0 and (A1.GT=1 or (A1.GT=0 and A0.GT=1))). Our test for GT is that ablation causes the answer to change sign. The $MT$ calculation is very similar.
We found $A_{n}.MT$ and $A_{n}.GT$ subtasks in all mixed models. Usually (but not always) both subtasks are calculated by the same node. (Table 2 shows an example.) As both the $MT$ and GT approaches are valid, models can learn valid algorithm sub-task implementations that differ per answer digit.
We further posit that a node that implements say A2.SA, A2.MD and A2.ND does not know whether it is dealing with a S, M or N question, so it outputs all three possible answers to the residual stream. Another node must calculate the S, M or N distinction likely by attending to the question operator (OPR) and the answer sign (SGN). We define attention sub-tasks OPR and SGN for these calculations. Table 2 shows that for each answer digit one attention head attends to both OPR and SGN. We believe these heads transfer sufficient data to the residual stream to allow the MLP layer(s) to calculate the question class, and so select the appropriate output from the polysemantic A2.SA/MD/ND node.
4.8 Mixed model summary
The 7 mixed experimental models achieve five or six nines accuracy and contain the same sub-tasks, implying a common algorithm. However, without fully understanding this algorithm, we can’t confirm these models as known-good.
5 Conclusion
We successfully trained and verified known-good 5-, 6- and 10-digit, 2-layer, 3-head addition models implementing the same algorithm with minor variations.
We demonstrated component reuse by integrating an existing addition model into a larger “mixed" model for both addition and subtraction, achieving six nines accuracy. This integration helped us understand the mixed model’s algorithm. The mixed model reuses most inserted addition nodes, upgrading many to become polysemantic - performing both addition and subtraction subtasks simultaneously.
Our work supports (Michaud et al., 2023)’s assertion that many natural prediction problems can be broken into finite "quanta" computations essential for loss minimization. It also aligns with the idea that understanding a network’s functionality involves identifying and comprehending its sub-quanta.
6 Limitations and Future Work
Our research suggests several future directions. A key goal is to complete reverse-engineering the mixed model’s algorithm, understanding polysemantic nodes and OPR+SGN node output representations in the residual stream, and how MLP layers use this to predict answer digits accurately.
We aim to develop a comprehensive, known-good model for n-digit addition, subtraction, and multiplication. Our approach also could be applied to create known-good models in logical reasoning and planning.
Using smaller models to enhance LLM trustworthiness and capabilities is an active research area, including model composition (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and inserting accurate models into untrained ones. Research on “spare” neurons in LLMs ((Voita et al., 2023; Hu et al., 2021)) suggests potential for small-scale modifications to fix erroneous circuits. Further exploration of using known-good models to improve LLMs is warranted.
Developing methods to incorporate compact known-good models into LLMs could democratize AI Safety research, allowing small teams to focus on specific areas and create quality components to improve LLMs.
Our automated framework for discovering algorithm sub-tasks in models accelerated our research and can test whether known-good components retain functionality when integrated into LLMs. While some aspects are specific to our math models, much is generic and reusable. Extending the framework to integrate output from complementary tools (e.g. ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights.
We progressed on a declarative language to describe algorithms in terms of necessary sub-tasks and a framework to test these descriptions against specific models. Further work is needed to fully realize its potential.
7 Impact Statement
Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications. A deeper understanding of generative AI has dual usage. While the potential for misuse exists, we discourage it. The knowledge gained can be harnessed to safeguard systems, ensuring they operate as intended. It is our sincere hope that this research will be directed towards the greater good, enriching our society and preventing detrimental effects. We encourage responsible use of AI, aligning with ethical guidelines.
Appendix A Appendix: Terminology
These terms and abbreviations are used in this paper and the associated Colabs and python code:
  • Pn : Model (input or output) token position. Zero-based. e.g. P18, P18L1H0
  • Ln : Model layer n. Zero-based. e.g. P18L1H2
  • Hn : Attention head n. Zero-based. e.g. P18L1H2
  • Mn : MLP neuron n. Zero-based
  • PnLnHn : Location / name of a single attention head, at a specified layer, at a specific token position
  • PnLnMn : Location / name of a single MLP neuron, at a specified layer, at a specific token position
  • D : First number of the pair question numbers
  • Dn : nth numeric token in the first question number. Zero-based. D0 is the units value
  • D’ : Second number of the pair question numbers
  • D’n : nth token in the second question number. Zero-based. D0 is the units value
  • A : Answer to the question (including answer sign)
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+" or "-" answer sign
  • S : Prefix for Addition. Think S for Sum. Aka ADD.
  • SA : Basic Add. An addition sub-task. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition sub-task. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
  • SS : Make Sum 9. An addition sub-task. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
  • ST : TriCase. An addition sub-task. Refer paper 2 for details
  • ST8, ST9, ST10 : Outputs of the ST TriCase sub-task.
  • M : Prefix for Subtraction with a positive answer. Think M for Minus. Aka SUB
  • : Basic Difference. A subtraction sub-task. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • : Borrow One. A positive-answer subtraction sub-task. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • MZ : Make Zero. A positive-answer subtraction sub-task. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • MT : TriCase. A positive-answer subtraction sub-task.
  • MT1, MT0, MT-1 : Outputs of the MT TriCase sub-task.
  • N : Prefix for Subtraction with a negative answer. Think N for Negative. Aka NEG
  • ND : Basic Difference. A negative-answer subtraction sub-task. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction sub-task. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • NZ : Make Zero. A negative-answer subtraction sub-task. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • NT : TriCase. A negative-answer subtraction sub-task.
  • GT : Greater Than. A (positive-answer or negative-answer) subtraction sub-task. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
  • OPR : Operator. A sub-task that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
  • SGN : Sign. A sub-task that attends to the first answer token, which is + or -
  • PCA : Principal Component Analysis
  • EVR : Explained Variance Ratio. In PCA, EVR represents the percentage of variance explained by each of the selected components.
Appendix B Appendix: Model Configuration
Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook. The Colab runs on a T4 GPU. Training takes up to 35 mins. The key parameters (and their common configurations) are:
  • n_layers = 1, 2 or 3: Number of layers.
  • n_heads = 3 or 4: Number of attention heads.
  • n_digits = 5, 6 or 10: Number of digits in the question.
Each digit is represented as a separate token. (Liu and Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers". The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation. Depending on the configuration, each training run processes 1 to 4 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
Addition and subtraction include rare edge cases. For example, the SS cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare. The data generator was enhanced to increase the frequency of all known edges cases. This lead to lower model loss.
Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
Appendix C Appendix: Model Loss
The model defaults to batch size = 64, learning rate = 0.00008 and weight decay = 0.1. The loss function is simple:
  • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
  • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
In our experimental models, the number of digits in the question varies from 5 to 10, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tab. 6.
Appendix D Appendix: Addition Model Shape
While we wanted a very low loss addition model, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model. Here are the things we tried to reduce loss that didn’t work:
  • Increasing the frequency of hard (cascading SS ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
  • Increasing the number of attention heads from 3 to 4 or 5 (while still using 1 layer) to provide more computing power.
  • Changing the question format from “12345+22222=” to “12345+22222equals” giving the model more prediction steps after the question is revealed before it needs to state the first answer digit.
  • With n_layers = 1 increasing the number of attention heads from 3 to 4.
  • Changing the n_layers to 2 and n_heads to 2.
The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads.
Appendix E Appendix: Experimental models
Twenty-one experiment models were trained and analyzed (refer Tab. 6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke/VerifiedArithmetic
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
Addition models
5 1 3 30K 372001 9.4e-2 12621 N/A 15 6
5 2 3 15K 372001 1.6e-8 0 N/A 30 16
5 2 3 40K 372001 2.0e-9 0 N/A 22 15
6 2 3 15K 372001 1.7e-8 2 N/A 31 17
6 2 3 20K 173289 1.5e-8 0 N/A 28 17
6 2 3 20K 572091 7.0e-9 0 N/A 35 17
6 2 3 40K 372001 2.0e-9 0 N/A 29 17
10 2 3 40K 572091 7.0e-9 0 N/A 44 28
Subtraction models
6 2 3 30K 372001 5.8e-6 N/A 0 40 21
10 2 3 75K 173289 2.0e-3 N/A 6672 101 37
Mixed models
6 3 4 40K 372001 5.0e-9 1 0 54 26
10 3 4 75K 173289 1.1e-6 2 295 143 53
Mixed models initialized with addition model
6 2 3 40K 572091 2.4e-8 0 5 57 21
6 3 3 40K 572091 1.8e-8 0 3 70 35
6 3 3 80K 572091 1.6e-8 0 3 75 35
6 3 4 40K 372001 8.0e-9 0 0 72 26
6 3 4 40K 173289 1.4e-8 3 2 60 29
6 3 4 50K 572091 2.9e-8 0 4 79 29
10 3 3 50K 572091 6.3e-7 6 7 90 45
Mixed models initialized with add model. Reset useful heads every 100 steps
6 4 4 40K 372001 1.7e-8 3 8 51 30
Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
6 4 3 40K 372001 3.0e-4 17 3120 115 53
Table 6: Main experimental models studied. The number of addition and subtraction failures per million questions is shown. The best 5-, 6- and 10-digit models are bolded.
For each model the ’VerifiedArithmeticTrain’ Colab notebook generates two files:
  • A “XXXXXX.pth" file containing the model weights
  • A “XXXXXX_train.json" file containing configuration information and training loss data
While, for each model the ’VerifiedArithmeticAnalysis’ Colab notebook generates two more files:
  • A “XXXXXX_behavior.json" file containing generic “behavior" facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
  • A “XXXXXX_maths.json" file containing “maths-specific" facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC sub-task.
Appendix F Appendix: TriAdd Implementation
TriAdd transfers data from $A_{n-1}$to $A_{n}$by integrating the values of $A_{n-1}.ST$ and $A_{n}.ST$ This function can be represented as nine bigram mappings with three possible outputs. (Refer Tab.7.)
$A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
$A_{n-1}.ST$ = ST8 = ST9 = ST10
ST8 ST8 ST9 ST10
ST9 ST8 ST9 ST10
 ST10 ST8 * ST10 ST10
Table 7: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate. The result could be ST8 or ST9 but importantly it can not be ST10. We choose to use ST8 in our definition, but ST9 would work just as well.
Appendix G Appendix: Complexity
To analyze question difficulty, we categorized addition questions by the complexity of the computation required to solve the question, as shown in Tab. 8. The categories are arranged according to the number of digits that a carry bit has to cascade through.
Name Contains Example Freq
S0 SA 11111+12345=23456 $\sim$5%
S1 SA,SC 11111+9=22230 $\sim$21%
S2 SA,SCx2 11111+89=22300 $\sim$34%
S3 SA,SCx3 11111+889=23000 $\sim$28%
S4 SA,SCx4 11111+8889=30000 $\sim$11%
S5 SA,SCx5 11111+88889=100000 $\sim$2%
Table 8: We categorise addition questions into non-overlapping “calculation complexity" quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
Appendix H Appendix: Addition Hypothesis 1
Given the 2-layer attention pattern’s similarity to 1-layer attention pattern, and the above evidence, our first (incorrect) hypothesis was that the 2-layer algorithm:
  • Is based on the same SA, SC and SS operations as the 1-layer.
  • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
  • The long double staircase still finalises each answer digit’s calculation.
  • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
If this is correct then the 2-layer algorithm successfully completes these calculations:
  • A0 = A0.SA
  • A1 = A1.SA + A0.SC
  • A2 = A2.SA + (A1.SC or (A1.SS & A0.SC))
  • A3 = A3.SA + (A2.SC or (A2.SS & A1.SC) or (A2.SS & A1.SS & A0.SC))
  • A4 = A4.SA + (A3.SC or (A3.SS & A2.SC) or (A3.SS & A2.SS & A1.SC) or (A3.SS & A2.SS & A1.SS & A0.SC))
  • A5 = A4.SC or (A4.SS & A3.SC) or (A4.SS & A3.SS & A2.SC) or (A4.SS & A3.SS & A2.SS & A1.SC) or (A4.SS & A3.SS & A2.SS & A1.SS & A0.SC)
Our intuition is that there are not enough useful nodes in positions 8 to 11 to complete the A5 calculation this way. So we abandoned this hypothesis.
Appendix I Appendix: Addition Hypothesis 2
Our second (incorrect) hypothesis was that the 2-layer algorithm has a more compact data representation, so it can pack more calculations into each node, allowing it to accurately predict A5 in step 11.
We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $A_{n}$.T, where T stands for “token addition”:
  • $A_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
The $A_{n}$.T operation does not understand mathematical addition. Tab. 9 shows how the model implements the T operator as a bigram mapping.
$D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
0 0 1 4 5 8 9
1 1 2 5 6 9 10
4 4 5 8 9 12 13
5 5 6 9 10 13 14
8 8 9 12 13 16 17
9 9 10 13 14 17 18
Table 9: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
$A_{n}$.T is a compact way to store data. Tab. 10 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
$A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
0 0 0 0
1 1 0 0
8 8 0 0
9 9 0 1
10 0 1 0
17 7 1 0
18 8 1 0
Table 10: Converting a $A_{n}$.T value into a SA , SC or SS value.
Our notation shorthand for one-digit-accuracy these “conversion" bigram mappings is:
  • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
  • $A_{n}$.SC = ($A_{n}$.T // 10) where // is the integer division operator
  • $A_{n}$.SS= ($A_{n}$.T == 9) where == is the equality operator
The A0.T value is accurate. But the other $A_{n}$.T values are not accurate because each is constrained to information from just one digit. We define another more accurate operator $A_{n}$.T2 that has “two-digit accuracy”. $A_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
  • $A_{n}$.T2 = $A_{n}$.T + $A_{n-1}$.SC
$A_{n}$.T2 is more accurate than $A_{n}$.T. The $A_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+CarryOne). Tab. 11 show how the model can implement the T2 operator as a mapping.
$A_{n}$.T $A_{n}$.T2 if $A_{n-1}$.SC==0 $A_{n}$.T2 if $A_{n-1}$.SC==1
0 0 1
1 1 2
9 9 10
10 10 11
17 17 18
18 18 19
Table 11: Calculating $A_{n}$.T2 from $A_{n}$.T and $A_{n-1}$.T
Following this pattern, we define operators $A_{n}$.T3, $A_{n}$.T4 and $A_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
  • $A_{n}$.T3 = $A_{n}$.T + ( $A_{n-1}$.T2 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC2
  • $A_{n}$.T4 = $A_{n}$.T + ( $A_{n-1}$.T3 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC3
  • $A_{n}$.T5 = $A_{n}$.T + ( $A_{n-1}$.T4 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC4
The value A4.T5 is accurate as it integrates SC and cascading SS data all the way back to and including A0.T. The values A1.T2, A2.T3, A3.T4 are also all accurate. If the model knows these values it can calculate answer digits accurately:
  • A1 = A1.T2 % 10
  • A2 = A2.T3 % 10
  • A3 = A3.T4 % 10
  • A4 = A4.T5 % 10
  • A5 = A4.T5 // 10
In this hypothesis, all the answer digits are accurately calculated using the nodes in positions 8 to 11. This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
Experimenation shows the model does not implement this hypothesis. It retains the long staircase SA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
  • Hypothesis 2 is too compact. The model is not optimising for compactness. The long staircase is discovered early in training, and it works for simple questions. Once the overall algorithm gives low loss consistently it stops optimising.
  • Hypothesis 2 accurately predicts all answer digits in step 11 - reflecting the authors (human) values for good code. The model is not motivated to do this. It just needs to accurately predict A5 as 1 or 0 in step 11 and A4 in step 12 - nothing more.
We abandoned this hypothesis.
Appendix J Appendix: Addition Hypothesis 3
The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Some of the experiments and mappings were:
  • Ablation experiments show that the A5 value is accurately calculated in prediction step 11 using 5 attention heads and 5 MLP layers. The pseudo-code accurately calculates A5 while constraining itself to this many steps.
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Table 1). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $A_{n}.ST$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 8).
  • The pseudo-code has 4 instances where compound functions using TriCase and TriAdd to generate tri-state outputs. PCA of the corresponding nodes (P11.L0.H1, P12.L0.H1 and P13.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 8).
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § K for details.
  • The pseudo code includes calculations like A1.ST which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of A1.ST to both nodes.
  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output (see Figure 9). The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state. We cannot see which. The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 9). The pseudo-code shows a bi-state calculation A0.SC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case. The model needs A1.ST2 for A2 accuracy. The model could simply reuse the accurate A1.ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST. That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values". This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to six nines accuracy.
Figure 5: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells. Each cell shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc. The model only uses nodes in nine token positions.
Figure 6: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P8 to P16 (i.e. tokens D’2 to A1). Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 7: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model. At token position P11 the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P11.
Figure 8: For 5-digit addition, for these attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA data and TriCase quanta are both tri-state and strongly correlated.
Figure 9: For 5-digit addition, for these attention heads, for exactly 1 digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA and TriCase data are strongly correlated, but the PCA data has 2 states.
Appendix K Appendix: Addition Interchange Interventions
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions" (aka activation patching) experiments were performed on the model, where
  • A particular claim about an attention head has selected for testing.
  • The model predicted answers for sample test questions, and the attention head activations were recorded (stored).
  • The model then predicted answers for more questions, but this time we intervened during the prediction to override the selected attention head activations with the activations from the previous run.
Using this approach we obtained the findings in Tab. 12.
Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
P8.L0.H1 and MLP A2.ST = TriCase( D2, D’2 ) impacting A4 and A5 accuracy Based on D2 and D’2. Triggers on a A2 carry value. Provides carry bit used in A5 and A4 calculation.
P9.L0.H1 and MLP A1.ST = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy Based on D1 and D’1. Triggers on a A1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
P10.L0.H1 and MLP A1.ST2 = TriAdd( A1.ST, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy Based on D0 and D’0. Triggers on a A0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
P11.L0.H1 and MLP A3.ST4 = TriAdd( TriCase( D3, D’3 ), TriAdd( A2.ST, A1.ST2)) impacting A5 accuracy Based on D3 and D’3. Triggers on a A3 carry value. Provides carry bit used in A5 calculations.
P11.L0.H2 and MLP A4.ST = TriCase(D4, D’4) impacting A5 accuracy A4 Based on D4 and D’4. Triggers on a A4 carry value. Provides carry bit used in A5 calculation.
P12.L0.H0+H2 and MLP A4.SA = (D4 + D’4) % 10 impacting A4 accuracy A4 Sums D4 and D’4. Impacts A4.
P13.L0.H0+H2 and MLP A3.SA = (D3 + D’3) % 10 impacting A3 accuracy Sums D3 and D’3. Impacts A3.
P14.L0.H0+H2 and MLP A2.SA = (D2 + D’2) % 10 impacting A2 accuracy Sums D2 and D’2. Impacts A2.
P14.L0.H1 and MLP (D1 + D’1) / 10 + P10.A1.ST2 info impacting A2 accuracy Calculates P10.A1.ST1 but add P10.A1.ST2 info when A1.ST != A1.ST2. Impacts A2
P15.L0.H0+H2 and MLP A1.SA = (D1 + D’1) % 10 impacting A1 accuracy Sums D1 and D’1. Impacts A1.
P15.L0.H1 and MLP A0.SC = (D0 + D’0) / 10 impacting A1 accuracy Triggers when D0 + D’0 >10. Impacts A1 digit by 1
P16.L0.H0+H2 and MLP A0.SA = (D0 + D’0) % 10 impacting A0 accuracy Sums D0 and D’0. Impacts A0.
Table 12: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head in a sample mixed model. Experimental results are consistent with hypothesis 3 for all nodes.
+ve Sub Contains -ve Sub Contains Like
M0 MD N/A N/A S0
M1 MD,MB N1 ND,NB S1
M2 MD,MBx2 N2 ND,NBx2 S2
M3 MD,MBx3 N3 ND,NBx3 S3
M4 MD,MBx4 N4 ND,NBx4 S4
Table 13: We define “positive-answer subtraction" and “negative-answer subtraction" calculation complexity quanta that parallel the addition quanta.
Appendix L Appendix: N-Digit Addition
The addition models perform addition accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
Figure 10: This map shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc.
Figure 11: This map shows the % of questions that fail when we ablate each node in the 6-digit 2-layer 3-head addition model. The model only uses nodes in token positions P11 to P20. Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not useful.
Some notes about the models:
  • The models selected different attention heads in the early positions to use to do the same logical calculations.
  • Some models use 2 attention heads per digit to do the SA calculation, whereas some models only uses one (and so are more compact).
  • The PCA trigrams have difference appearances in different models (but the same interpretable clusters). Refer Figures 8
  • Per answer digit, some models use the SC calculation, whereas some models optimize it out and rely solely on the $ST$ value (and so are more compact).
Appendix M Appendix: Mixed Model Initialization
We experimented with three approaches to re-using the trained addition model in the “mixed" (addition and subtraction) model:
  • Initialize Only: Initialize the untrained mixed model with the addition model weights before training begins.
  • Freeze Attention: As per “Initialize Only", but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
  • Freeze All: As per “Initialize Only", but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
Our intuition was that “Initialize Only" would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention" and “Freeze All" approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
After experimentation we found that the “Initialize Only" approach was the only one that quickly trained to be able to do both addition and subtraction accurately. We concluded that the other two methods constrainws the model’s ability to learn new algorithms too much.
We also experimented with “where" in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model. That is, do we initialize the first 2 layers or the last 2 layers of the mixed model? Also do we initialize the first 3 attention heads or the last 3 attention heads of the mixed model? Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
Appendix N Appendix: N-Digit Subtraction
The mixed models perform addition and subtraction accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
Figure 12: This map of a sample 6-digit mixed model shows the 98 nodes used to predict answers to addition (S), positive-answer subtraction (M) and/or negative-answer subtraction (N) questions. Before training the mixed model, 48 nodes were initalized pre-training with a smaller addition model’s weights. These are have a red border. During mixed model training, 39 of 48 of the initalized monosemantic nodes were generalized (become poly-semantic) and now help predict two or three question classes.
Figure 13: This map shows the simplest complexity quanta S0, S1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model when doing addition questions.
Figure 14: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers.
Some notes about the mixed models:
  • All the notes about the addition model (above) also apply to the mixed model.
  • The model contains a new sub-task that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8" or “9" then the first answer token is more likely to be a “+" (and not a “-"). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.