### v9

# Arithmetic in Transformers Explained

Abstract
While recent work has shown transformers can learn addition, previous models exhibit poor prediction accuracy and are limited to small numbers. Furthermore, the relationship between single-task and multitask arithmetic capabilities remains unexplored. In this work, we analyze 44 autoregressive transformer models trained on addition, subtraction, or both. These include 16 addition-only models, 2 subtraction-only models, 8 “mixed” models trained to perform addition and subtraction, and 14 mixed models initialized with parameters from an addition-only model. The models span 5- to 15-digit questions, 2 to 4 attention heads, and 2 to 3 layers. We show that the addition models converge on a common logical algorithm, with most models achieving >99.999% prediction accuracy. We provide a detailed mechanistic explanation of how this algorithm is implemented within the network architecture. Subtraction-only models have lower accuracy. With the initialized mixed models, through parameter transfer experiments, we explore how multitask learning dynamics evolve, revealing that some features originally specialized for addition become polysemantic, serving both operations, and boosting subtraction accuracy. We explain the mixed algorithm mechanically. Finally, we introduce a reusable library of mechanistic interpretability tools to define, locate, and visualize these algorithmic circuits across multiple models.
Machine Learning, ICML
11footnotetext: Corresponding author: philipquirkenz@gmail.com
1 Introduction
Figure 1: An overview of our methodology: (1) We trained an accurate 5-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need to understand their behavior. Hendrycks & Mazeika (2022) argue that making these models interpretable is key to their safe use.
Previous work by (Quirke & Barez, 2024) demonstrated that transformers can learn addition on small numbers, but their models had poor accuracy and were limited to a single operation. Understanding how transformers can overcome these limitations to perform more accurate mathematical computations across multiple operations is crucial for several reasons. First, it provides insights into how neural networks can learn and combine multiple precise algorithms, a key capability for more complex reasoning tasks. Second, studying how models transition from single-task to multitask competence while maintaining accuracy reveals important principles about knowledge transfer and representation sharing in neural networks.
Understanding how models arrive at their predictions is critical for diagnosing errors, enhancing safety, and gaining fundamental insights into modern neural networks. Mechanistic interpretability provides a pathway toward this understanding by identifying the “circuits” responsible for particular computations within a network’s layers and attention heads.
In this work, we use multidigit arithmetic as a revealing testbed for mechanistic interpretability. We train and analyze 44 autoregressive transformer models on addition, subtraction, or both, covering question lengths of 5 to 15 digits and varying in their number of layers (2–3) and attention heads (2–4). Our experiments encompass three categories of models: addition-only, subtraction-only, and “mixed” models trained to handle both operations. Through a systematic investigation of these networks, we show that addition-only models converge on a highly accurate, shared algorithm, yielding near-perfect performance. We mechanistically explain the algorithm. Subtraction-only models, in contrast, exhibit lower accuracy.
To explore how multitask learning reshapes these internal algorithms, we then conduct parameter transfer experiments using three distinct approaches, each involving the initialization of mixed models with parameters from addition-only networks. Interestingly, one of these approaches boosts the subtraction accuracy beyond what is achieved by mixed models trained from scratch, indicating that features specialized for addition can be effectively repurposed for subtraction. We provide a detailed mechanistic explanation of the mixed algorithm, illustrating how certain addition circuits become “polysemantic” under multitask learning, driving behavior for both operations.
A further contribution of our study is the development of a reusable library of mechanistic interpretability tools. This library enables researchers to define and search for algorithmic features, visualize network structures and feature interactions, and conduct cross-model comparisons efficiently.
Overall, our results demonstrate the power of mechanistic interpretability in shedding light on how transformers learn and reuse algorithmic circuits across related tasks, paving the way for deeper understanding and more reliable deployment of large language models in diverse domains.
2 Related Work
Mechanistic interpretability aims to reverse engineer neural networks to find interpretable algorithms that are implemented in a model’s weights (Olah et al., 2020). Mathematical frameworks (Elhage et al., 2021a) explain how transformer attention heads can work with each other to implement complex algorithms.
Causal Scrubbing (Jenner et al., 2023) recommends explaining a model algorithm by documenting a low-level computation graph, mapping from the graph to the model nodes that implement the computation, and performing experimentation verification. Investigative techniques such as ablation interventions, activation unembeddings (nostalgebraist, 2020), and sparse autoencoders (Nanda, 2023; Cunningham et al., 2023), underpinned by the more theoretical frameworks (Elhage et al., 2021b; Geva et al., 2022), provide tools to help confirm a mapping.
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems. It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu & Low, 2023).
Studying Toy Models for Arithmetic. Applying mechanistic interpretability to toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023). Quirke & Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on edge cases (e.g. “77778+22222=100000” where a “carry 1” cascades through 4 digits), incorrectly predicting 1% of questions.
Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta) (Michaud et al., 2023). Models must learn these quanta to reduce loss. Understanding a network reduces to enumerating its quanta. Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
3 Methodology
Transformer models learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implements this approach.
3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. We define the first number (or addend) as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer (or sum) as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. Figure 2 shows an illustrative example.
Figure 2: For 5-digit addition, our model has 12 input (question) and 6 output (answer) token positions. We name the question tokens D4, …, D0, and D’4, …, D’0 and the answer tokens A5, …, A0. For n-digits, we use the terms $D_{n}$, $D^{\prime}_{n}$and $A_{n}$.
We adopt the framework from Quirke & Barez (2024) for our model’s addition process, where they break down the task of multi-digit addition into an addition subtask and a carry subtask. Firstly, the “Simple Addition” subtask $A_{n}.SA$ naively calculates the sum of corresponding digit pairs in the addends:
$A_{n}.SA=(D_{n}+D^{\prime}_{n})\mod 10$ (1)
The “Simple Carry” subtask $A_{n}.SC$ then determines whether the addition between the two digits creates a carry bit:
$A_{n}.SC=\begin{cases}1&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ 0&\text{otherwise}.\end{cases}$ (2)
By combining these two subtasks, the model and framework from Quirke & Barez (2024) is capable of handling simple addition and carry bits generated directly from digit pair addition. However, it struggles with problems with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
Consider “00144+00056=000210”. Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position. A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110”. The Quirke & Barez (2024) 1-layer model could cascade carry bits across two digits, but not three or more.
3.2 Extending the Mathematical Framework
To solve $n$-digit addition with high accuracy, we need to extend the basic framework to handle carry bits that cascade through multiple digits. Consider the problem ”44444+55556=100000”. For an autoregressive model to correctly predict the first answer digit as ”1”, it needs to track how the carry bit cascades from the rightmost digit all the way to the leftmost position in a single forward pass. This is particularly challenging for an autoregressive model that predicts answer tokens from left to right.
To solve this, we extend the mathematical framework with two key subtasks.
First, the digit-level subtask ”Sum Type” $A_{n}.ST$ analyzes each pair of digits and classifies their sum as either definitely causing a carry, possibly causing a carry, or definitely not causing a carry. This is implemented by a TriCase operation that maps each digit pair ($D_{n}$ and $D^{\prime}_{n}$), TriCase into one of three possible states:
$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10\text{ (definite carry)}\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9\text{ (possible carry)}\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8\text{ (no carry)}\\ \end{cases}\end{split}$ (3)
The key insight is that while some digit pairs definitely generate a carry (like 5+6), and others definitely don’t (like 2+3), pairs that sum to 9 (like 4+5) may or may not generate a carry depending on whether there’s an incoming carry bit from the next lower digit.
Then, we introduce the ”Sum Validation” subtask $A_{n}.SV$ that processes these carry states sequentially to determine whether any ”possible carries” become ”definite carries”. This subtask is implemented using a TriAdd operation that examines each digit’s sum type along with the sum type of the next lower digit:
$\begin{split}A_{n}.SV&=\underbrace{\text{TriAdd}}_{(A_{n}.ST,A_{n-1}.ST)}=\\ &\begin{cases}ST10&\text{if }A_{n}.ST=ST10\text{ or }\\ &\quad(A_{n}.ST=ST9\text{ and }\\ &\quad A_{n-1}.ST=ST10),\\ ST8&\text{otherwise.}\end{cases}\end{split}$ (4)
Revisiting the problem ”44444+55556=100000”, the model can accurately determine whether the first answer digit A5 should be 1 or 0 by examining all digit pairs in sequence. This chain of operations uses TriAdd to process the Sum Types from least to most significant digits:
$A1.SV$ $=\text{TriAdd}(A1.ST,A0.ST)$
$A2.SV$ $=\text{TriAdd}(A2.ST,A1.SV)$
$A3.SV$ $=\text{TriAdd}(A3.ST,A2.SV)$
$A4.SV$ $=\text{TriAdd}(A4.ST,A3.SV)$
$A5$ $=(A4.SV=ST10)$
Most importantly, by using this cascading calculation to calculating A5, the model also simulatenously calculates the accurate carry bits for each answer digit. The carry bits are simply the SV values we computed along the way—we already have A3.SV for A4’s carry bit, A2.SV for A3’s carry bit, and so on.
Within our extended framework, the model only requires the subtasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform $n$-digit addition accurately. Some models may also use the redundant $A_{n}$.SC subtask.
Figure 3 shows how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke & Barez (2024) algorithm, but contains an additional circuit (the shaded boxes) to calculate “cascading carry one” data to solve $n$-digit addition with >99.999% accuracy.
Figure 3: To predict with >99.999% accuracy, the addition algorithm first calculates “carry one” values $(A_{n}.ST)$, combining them into “cascading carry one” values $(A_{n}.SV).$ At the “+” token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0. The other answer digits are calculated by combining “base add” $(A_{n}.BA)$ and “carry one” $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
3.3 Techniques
To undetsand how the model implements arithmetic operations, we analyzed both attention heads and MLP layers at each token position. We define a node as either an attention head or MLP layer at a specific token position (e.g., the first attention head in the first layer processing the third digit). We employed four techniques to understand each node’s function:
  1. 1.
    Intervention Ablation. To determine the role of a node’s output, we replace its output with the mean vector of its outputs across a batch, then measured how this impacted downstream predictions. For each subtask, we developed automated intervention ablation tests across $n$-digits to verify whether a node implemented that subtask.
  2. 2.
    Attention Patterns. For each attention head node, we examined which input tokens received significant attention (defined as post-softmax attention score $>0.01$). This revealed which parts of the input each attention head considered relevant.
  3. 3.
    Principal Component Analysis (PCA). We applied PCA to node outputs, especially when investigating attention heads that our framework suggests might have descrete states (e.g. tri-state or bi-state outputs).
  4. 4.
    Question Complexity. We categorized questions by computational complexity (App. G). For addition, categories S0-S4 reflect how many sequential digits a “carry one” operation cascades through. For subtraction, categories M0-M4 reflect a similar cascading of ”borrow one” operations. We analyzed which nodes were necessary for correct predictions in each category.
4 Experiments
4.1 Training a Five-Digit Addition Model
Previous work by Quirke & Barez (2024) achieved $\sim$ 99% accuracy on 5-digit addition using a 1-layer addition model. To improve upon this, we systematically explored different model architectures and found that a transformer with 2 layers and 3 attention heads per layer was the minimal configuration needed to achieve >99.999% accuracy (see App. D for our exploration of alternative architectures, and App. B for model configuration details.). This 2-layer architecture doubles the model’s computational capacity and enables cross-layer attention composition for more sophiscated algorithms (Elhage et al., 2021b).
We trained our model to perform 5-digit addition using an architecture with 2 layers and 3 attention heads per layer. The model’s vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /. We used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1.
To ensure robust performance on edge cases, we trained on an infinite dataset that was deliberately enriched with rare numerical combinations. The loss function was defined as the mean negative log likelihood across all output tokens. After 30,000 training batches, the model achieved a final training loss of approximately $2.3\times 10^{-8}$, and subsequent testing confirmed accuracy exceeding 99.999%. For comprehensive training details and results, see App. C and Tab. 6.
Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the ”=” token, each node often calculates data used to predict multiple answer digits. After the ”=” token, all nodes in a given token position are used to predict a single answer digit.
(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
D’4 D’1 D’0 = + A5 A4 A3 A2 A1
L0H0 A4 A2 A1
L0H1 A5 A5..3 A5 A3 A0
L0H2 A5 A5..1
L0MLP A5..2 A4 A3 A2 A1 A0
L1H0
L1H2 A4 A3 A2
L1MLP A1 A0
4.2 Investigating Five-Digit Addition
Through ablation experiments, we found that the model relies on just 36 nodes distributed across nine token positions (Figs 7 and 8). We analyzed how these nodes affect computation complexity and answer accuracy through ablation studies (see Figs 6 and Table 1).
Core Algorithm Discovery. The model’s most challenging task is predicting A5, the first answer digit, at position P11. This digit, which can only be 0 or 1, depends on potential carry-bit cascades through multiple positions (e.g., in 55555+44445=100000). After two unsuccessful algorithm hypotheses (detailed in App.H andI), we discovered the model’s fundamental mechanism: it uses a minimal representation of carry information through what we call the TriCase quanta. The model computes $A_{n}.ST$ using bigram operations that map digit pairs to result tokens (e.g., ”6” + ”7” = ST10) in positions P8 to P11.
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention. Figure 3 diagrams the algorithm, with node details in App. J. The model adheres to all known constraints and achieves >99.999% accuracy using 36 nodes, as compared to the 1-layer model uses 21 nodes to achieve 99% accuracy.
Table 2: All addition models studied implement our addition algorithm. The algorithm SA, SC and $ST$ subtasks all exist for each digit and in appropriate token positions. For a sample model, this map shows the subtask locations. Interestingly, here each SA subtask is shared across two attention heads.
(P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’1 D’0 = + A6 A5 A4 A3 A2 A1
L0H0 A2.ST A3.ST A1.ST A4.ST A4.SC A3.SC A2.SC A1.SC A0.SC
L0H1 A1.ST A0.ST A5.SA A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST
4.3 Investigating n-digit Addition models
To test whether this algorithm generalises beyond 5-digit addition, we conducted a comprehensive study of larger addition models. We trained 15 2-layer transformer models to handle addition problems ranging from 5- to 15-digits. These models used varying random seeds and optimizers, and included a sign token in their output format (e.g. 111111+222222=+0333333).
Figure 4: The 5- to 13-digit 2-layer 3-head addition models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.6.
Analysis Framework. To systematically analyse these models, we first created a declarative method for specifying each subtask within the addition process. Each subtask definition includes (1) the input digits it processes, (2) the answer digit it affects, (3) the complexity level it influences, and (4) expected PCA patterns and ablation test results. Using this declaration, we identified the nodes performing these subtasks across all models.
Since our framework lays out the addition algorithm in terms of its required subtasks., we can also verify whether a given model implements this algorithm by checking for the presence and proper sequencing of all required subtasks. For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ subtasks for each question digit before the first answer token. Table 2 shows how an addition model fully implements all required subtasks in its nodes.
Using this framework, we found that all fifteen 2-layer addition models, regardless of their size, implemented the same core algorithm we identified in our 5-digit model. This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup where the digits are tokenized individually.
4.4 Training n-digit Mixed models
To explore how transformers can learn multiple arithmetic operations simultaneously, we conducted experiments with models capable of both addition and subtraction. We were particularly interested in whether knowledge from addition-specialized models could transfer to this expanded task.
We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads. Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model. During training, we used a curriculum of 80% subtraction and 20% addition examples, with deliberate enrichment of rare edge cases for both operations.
This initialization approach proved effective, with most models achieving high accuracy on both operations. We explored alternative training strategies, such as periodically ”freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps. However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance.
The initialized mixed experimental models achieve >99.999% accuracy and contain the same subtasks, implying a common algorithm.
Figure 5: The 5- to 8-digit initialized mixed models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.7.
4.5 Investigating n-digit Mixed models
The mixed models revealed fascinating patterns in how they handled different types of arithmetic problems. Our hypothesis was that the model handled three distinct question classes: addition, positive-answer subtraction (e.g., 10009-10000=+000009), and negative-answer subtraction (e.g., 10009-20000=-009991).
Through ablation studies, we found that many nodes specialized in handling specific question classes, while others became ”polysemantic,” capable of processing multiple classes simultaneously (More details in Tab.3). Most notably, the majority of nodes initialized from the addition model became polysemantic during training. Rather than developing entirely new circuits for subtraction, the models adapted their addition circuits to handle both operations.
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction subtasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
Used Inserted
Question class # % # %
All questions 96 48
Addition 61 64% 42 88%
Positive-answer sub 70 73% 40 83%
Negative-answer sub 53 55% 29 60%
Defining Subtraction Subtasks. Subtraction requires two distinct computational paths based on the relative magnitude of the numbers. When D $\geq$ D’, as in 325 - 123 = +202, we perform direct subtraction with borrowing. When D $<$ D’, as in 325 - 456 = -131, we must instead compute the magnitude of the difference and mark it negative.
Hence, this computational difference necessitates separate frameworks for positive-answer (M) and negative-answer (N) subtraction, each with parallel operations to addition’s Base Add ($SA$) and Carry One ($SC$), which we define in Tab. 4. We also define complexity measures for subtraction (see App. G and  N).
Table 4: We define 2 “positive-answer subtraction” and 2 “negative-answer subtraction” subtasks that parallel the addition subtasks, Base Add SA and Carry One SC .
Name +ve Sub -ve Sub Definition
Base Diff MD ND $D_{n}$- $D^{\prime}_{n}$10
Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
Just as addition must handle ”cascading carry” cases (e.g., 99999+1=100000), subtraction must handle ”cascading borrow” cases. Consider 100000-1=099999, where a borrow cascades through multiple zeros. To handle such cases efficiently, we define a three-state classification $MT$ that parallels addition’s $ST$ :
$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one. For N questions, the interpretation is the opposite.
Finally, paralleling the ”cascading carry one” $An.SV$ operation in addition, we define cascading borrow calculations $A_{n}.MV$ and $A_{n}.NV$ for positive and negative cases.
Evolution of Polysemantic Nodes. During training, the mixed model adapted the nodes it inherited from the addition model. Nodes that originally only performed Simple Addition ($SA$) developed the ability to simultaneously handle three distinct operations: addition ($SA$), positive-answer subtraction ($MD$), and negative-answer subtraction ($ND$). (Refer to Table 5.) We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs). The model similarly adapted nodes that originally handled carry bits ($SC$) to also process borrow operations for both positive ($MB$) and negative ($NB$) subtraction cases.
Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N. Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token. The MLP layers then select the output appropriate for the class. In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
L0H0 A4.MT A3.MT A4.ST A2.SC A1.SC A0.SC OPR
A4.GT A3.MT A4.MT A4.SC A3.SC A2.NB A1.MB A0.MB SGN
OPR A1.NB A0.NB
L0H1 A2.ST A1.ST A3.ST A0.ST A5.SA
A4.ST A2.MT A1.MT A3.GT A0.MT A5.MD
A2.GT A1.GT A0.GT A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST A5.SA A4.MD A3.MD A2.MD A1.MD A0.MD
OPR A5.MD A4.ND A3.ND A2.ND A1.ND A0.ND
SGN A5.ND
L0H3 OPR OPR OPR OPR OPR OPR
SGN SGN SGN SGN SGN SGN
4.6 Mixed model question class detection
Mixed models must identify whether each input belongs to one of three classes: addition (S), positive-answer subtraction (M), or negative-answer subtraction (N). For addition questions (marked by ”+”), classification is immediate. For subtraction questions (marked by ”-”), the model must determine whether D $\geq$ D’ to distinguish between M and N classes. This classification must occur by the ”=” token since it determines whether the first answer token will be ”+” or ”-”.
The model can determine D $\geq$ D’ in two ways. First, it can use the $A_{n}.MT$ values already computed for subtraction. Alternatively, it can use a dedicated comparison circuit based on a new subtask:
$A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
For a 4-digit number, this comparison cascades through the digits: D $\geq$ D’ is true if either $A3.GT=1$, or if $A3.GT=0$ and $A2.GT=1$, and so on through $A0.GT$. We verify this circuit through ablation tests that show its disruption causes answer signs to flip.
We found all mixed models incorporate the $A_{n}.MT$ and $A_{n}.GT$ subtasks to determine whether the answer is postiive or negative. Typically (but not always), the same node computes both subtasks (see Table 2). Since both the $MT$ and GT approaches can validly distinguish between positive and negative outcomes, the model sometimes chooses one approach for some digits and the other approach for others.
A single node can implement multiple arithmetic operations (e.g., $A2.SA$, $A2.MD$, $A2.ND$) but has no intrinsic way of knowing whether the current question is addition, positive subtraction, or negative subtraction. Hence, we hypothesise that it outputs all three possible results into the residual stream. A different node must then select the correct result by identifying the question’s operation and required sign.
In practice, we observe that one or more attention heads attend to the question operator token (OPR) and the initial answer sign token (SGN), encoding the necessary information for the downstream MLP layers to pick which of the three outputs is correct. As Table 2 shows, each answer digit typically has one attention head that focuses on both the OPR and SGN tokens, thereby enabling the MLP to filter out the irrelevant outputs and retain the correct one.
5 Conclusion
Our analysis of 44 transformer models demonstrates their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 13 digits. By investigating how models implement computations using attention heads and MLPs, we uncovered novel computational approaches distinct from traditional human calculation methods. Parameter transfer experiments revealed that features originally specialized for addition could be repurposed for subtraction, with many nodes becoming polysemantic.
Our research provides insights into how transformers learn precise algorithmic computations across related tasks, developing a reusable library of mechanistic interpretability tools that advances our understanding of neural network computational learning.
Impact Statement
Our work aims to explain the inner workings of transformer-based language models, which may have broad implications for a wide range of applications. A deeper understanding of generative AI has dual usage. While the potential for misuse exists, we discourage it. The knowledge gained can be harnessed to safeguard systems, ensuring they operate as intended. It is our sincere hope that this research will be directed towards the greater good, enriching our society and preventing detrimental effects. We encourage responsible use of AI, aligning with ethical guidelines.
Appendix A Appendix: Terminology
These terms and abbreviations are used in this paper and the associated Colabs and python code:
  • Pn : Model (input or output) token position. Zero-based. e.g. P18, P18L1H0
  • Ln : Model layer n. Zero-based. e.g. P18L1H2
  • Hn : Attention head n. Zero-based. e.g. P18L1H2
  • Mn : MLP neuron n. Zero-based
  • PnLnHn : Location / name of a single attention head, at a specified layer, at a specific token position
  • PnLnMn : Location / name of a single MLP neuron, at a specified layer, at a specific token position
  • D : First number of the pair question numbers
  • Dn : nth numeric token in the first question number. Zero-based. D0 is the units value
  • D’ : Second number of the pair question numbers
  • D’n : nth token in the second question number. Zero-based. D0 is the units value
  • A : Answer to the question (including answer sign)
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or ”-” answer sign
  • S : Prefix for Addition. Think S for Sum. Aka ADD.
  • SA : Basic Add. An addition subtask. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition subtask. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
  • SS : Make Sum 9. An addition subtask. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
  • ST : TriCase. An addition subtask. Refer paper 2 for details
  • ST8, ST9, ST10 : Outputs of the ST TriCase subtask.
  • M : Prefix for Subtraction with a positive answer. Think M for Minus. Aka SUB
  • : Basic Difference. A subtraction subtask. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • : Borrow One. A positive-answer subtraction subtask. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • MZ : Make Zero. A positive-answer subtraction subtask. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • MT : TriCase. A positive-answer subtraction subtask.
  • MT1, MT0, MT-1 : Outputs of the MT TriCase subtask.
  • N : Prefix for Subtraction with a negative answer. Think N for Negative. Aka NEG
  • ND : Basic Difference. A negative-answer subtraction subtask. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction subtask. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • NZ : Make Zero. A negative-answer subtraction subtask. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • NT : TriCase. A negative-answer subtraction subtask.
  • GT : Greater Than. A (positive-answer or negative-answer) subtraction subtask. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
  • OPR : Operator. A subtask that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
  • SGN : Sign. A subtask that attends to the first answer token, which is + or -
  • PCA : Principal Component Analysis
  • EVR : Explained Variance Ratio. In PCA, EVR represents the percentage of variance explained by each of the selected components.
Appendix B Appendix: Model Configuration
Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook. The Colab runs on a T4 GPU. Each training run takes up to 60 mins. The key parameters (and their common configurations) are:
  • n_layers = 1, 2 or 3: Number of layers.
  • n_heads = 3 or 4: Number of attention heads.
  • n_digits = 5, 6 or 10: Number of digits in the question.
Each digit is represented as a separate token. (Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”. The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation. Depending on the configuration, each training run processes 1 to 4 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
Addition and subtraction include rare edge cases. For example, the SS cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare. The data generator was enhanced to increase the frequency of all known edges cases. This lead to lower model loss.
Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
The Colabs will be made available on publication.
Appendix C Appendix: Model Loss
The model defaults to batch size = 64, learning rate = 0.00008 and weight decay = 0.1. The loss function is simple:
  • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
  • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
In our experimental models, the number of digits in the question varies from 5 to 13, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tabs. 6 and  7.
Appendix D Appendix: Addition Model Shape
While we wanted a very low loss addition model, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model. Here are the things we tried to reduce loss that didn’t work:
  • Increasing the frequency of hard (cascading SS ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
  • Increasing the number of attention heads from 3 to 4 or 5 (while still using 1 layer) to provide more computing power.
  • Changing the question format from “12345+22222=” to “12345+22222equals” giving the model more prediction steps after the question is revealed before it needs to state the first answer digit.
  • With n_layers = 1 increasing the number of attention heads from 3 to 4.
  • Changing the n_layers to 2 and n_heads to 2.
The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads.
Appendix E Appendix: Experimental models
Forty-four models were trained and analyzed (Refer Tabs. 6 and  7). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke in the quanta_maths_* collections to support further research in AI Safety.
Table 6: Addition-only models studied. The number of addition failures per million questions is shown. The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 7.
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Heads used MLPs used
5 1 3 30K 372001 9.4e-2 12621 15 6
5 2 3 15K 372001 1.6e-8 0 30 16
5 2 3 40K 372001 2.0e-9 0 22 15
6 2 3 15K 372001 1.7e-8 2 31 17
6 2 3 20K 173289 1.5e-8 0 28 17
6 2 3 20K 572091 7.0e-9 0 35 17
6 2 3 40K 372001 2.0e-9 0 29 17
7 2 3 45K 173289 3.0e-9 0 31 20
8 2 3 45K 173289 3.0e-9 0 35 22
9 2 3 45K 173289 3.0e-9 0 54 27
10 2 3 40K 572091 7.0e-9 0 44 28
11 2 3 50K 173289 8.0e-9 2 56 29
12 2 3 50K 173289 5.0e-9 3 50 33
13 2 3 50K 173289 6.3e-8 1 66 31
14 2 3 60K 173289 5.5e-6 199 68 35
15 2 3 80K 572091 8.6e-8 10 93 58
Table 7: Subtraction-only and mixed models studied. The number of addition and subtraction failures per million questions is shown. The number of useful attention heads at token position and useful MLP layers at token position are shown - summarizing the data shown in figures like Fig 7.
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
Subtraction models
6 2 3 30K 372001 5.8e-6 N/A 0 40 21
10 2 3 75K 173289 2.0e-3 N/A 6672 101 37
Mixed models
5 3 4 40K 372001 9.0e-9 0 0 45 21
6 3 4 40K 372001 5.0e-9 1 0 54 26
7 3 4 50K 372001 2.0e-8 2 6 108 40
8 3 4 60K 173289 4.7e-8 0 7 123 45
9 3 4 60K 173289 3.2e-7 1 33 140 46
10 3 4 75K 173289 1.1e-6 2 295 143 53
11 3 4 80K 572091 3.9e-8 0 13 138 50
12 3 4 85K 572091 1.7e-8 2 10 167 55
13 3 4 685K 572091 9.5e-6 399 4164 197 64
Mixed models initialized with addition model
5 2 3 40K 572091 2.5e-7 1 52 49 20
6 2 3 40K 572091 2.4e-8 0 5 57 21
6 3 3 40K 572091 1.8e-8 0 3 70 35
6 3 3 80K 572091 1.6e-8 0 3 75 35
6 3 4 40K 372001 8.0e-9 0 0 72 26
6 3 4 40K 173289 1.4e-8 3 2 60 29
6 3 4 50K 572091 2.9e-8 0 4 79 29
7 3 4 50K 572091 1.8e-8 4 1 104 38
8 3 4 70K 572091 4.3e-5 50 1196 116 42
9 3 4 70K 572091 5.4e-8 1 4 160 50
10 3 3 50K 572091 6.3e-7 6 7 90 45
11 3 4 75K 572091 5.9e-5 11066 1120 141 47
Mixed models initialized with add model. Reset useful heads every 100 steps
6 4 4 40K 372001 1.7e-8 3 8 51 30
Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
6 4 3 40K 372001 3.0e-4 17 3120 115 53
For each model the ’QuantaMathsTrain’ Colab notebook generates two files:
  • A “model.pth” file containing the model weights
  • A “training_loss.json” file containing configuration information and training loss data
While, for each model the ’QuantaMathsAnalysis’ Colab notebook generates two more files:
  • A “behaviors.json” file containing generic “behavior” facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
  • A “features.json” file containing maths-specific ”feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC subtask.
Appendix F Appendix: TriAdd Implementation
TriAdd transfers data from $A_{n-1}$to $A_{n}$by integrating the values of $A_{n-1}.ST$ and $A_{n}.ST$ This function can be represented as nine bigram mappings with three possible outputs. (Refer Tab.8.)
Table 8: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
$A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
$A_{n-1}.ST$ = ST8 = ST9 = ST10
ST8 ST8 ST9 ST10
ST9 ST8 ST9 ST10
 ST10 ST8 * ST10 ST10
Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate. The result could be ST8 or ST9 but importantly it can not be ST10. We choose to use ST8 in our definition, but ST9 would work just as well.
Appendix G Appendix: Complexity
To analyze question difficulty, we categorized addition questions by the complexity of the computation required to solve the question, as shown in Tab. 9. The categories are arranged according to the number of digits that a carry bit has to cascade through.
Table 9: We categorise addition questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
Name Contains Example Freq
S0 SA 11111+12345=23456 $\sim$5%
S1 SA,SC 11111+9=22230 $\sim$21%
S2 SA,SCx2 11111+89=22300 $\sim$34%
S3 SA,SCx3 11111+889=23000 $\sim$28%
S4 SA,SCx4 11111+8889=30000 $\sim$11%
S5 SA,SCx5 11111+88889=100000 $\sim$2%
Appendix H Appendix: Addition Hypothesis 1
Given the 2-layer attention pattern’s similarity to 1-layer attention pattern, and the above evidence, our first (incorrect) hypothesis was that the 2-layer algorithm:
  • Is based on the same SA, SC and SS operations as the 1-layer.
  • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
  • The long double staircase still finalizes each answer digit’s calculation.
  • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
If this is correct then the 2-layer algorithm successfully completes these calculations:
  • A0 = A0.SA
  • A1 = A1.SA + A0.SC
  • A2 = A2.SA + (A1.SC or (A1.SS & A0.SC))
  • A3 = A3.SA + (A2.SC or (A2.SS & A1.SC) or (A2.SS & A1.SS & A0.SC))
  • A4 = A4.SA + (A3.SC or (A3.SS & A2.SC) or (A3.SS & A2.SS & A1.SC) or (A3.SS & A2.SS & A1.SS & A0.SC))
  • A5 = A4.SC or (A4.SS & A3.SC) or (A4.SS & A3.SS & A2.SC) or (A4.SS & A3.SS & A2.SS & A1.SC) or (A4.SS & A3.SS & A2.SS & A1.SS & A0.SC)
Our intuition is that there are not enough useful nodes in positions 8 to 11 to complete the A5 calculation this way. So we abandoned this hypothesis.
Appendix I Appendix: Addition Hypothesis 2
Our second (incorrect) hypothesis was that the 2-layer algorithm has a more compact data representation, so it can pack more calculations into each node, allowing it to accurately predict A5 in step 11.
We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $A_{n}$.T, where T stands for “token addition”:
  • $A_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
The $A_{n}$.T operation does not understand mathematical addition. Tab. 10 shows how the model implements the T operator as a bigram mapping.
Table 10: Implementing the T operator as a bigram mapping from 2 input tokens to 1 result token.
$D_{n}$vs $D^{\prime}_{n}$ 0 1 4 5 8 9
0 0 1 4 5 8 9
1 1 2 5 6 9 10
4 4 5 8 9 12 13
5 5 6 9 10 13 14
8 8 9 12 13 16 17
9 9 10 13 14 17 18
$A_{n}$.T is a compact way to store data. Tab. 11 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
Table 11: Converting a $A_{n}$.T value into a SA , SC or SS value.
$A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
0 0 0 0
1 1 0 0
8 8 0 0
9 9 0 1
10 0 1 0
17 7 1 0
18 8 1 0
Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
  • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
  • $A_{n}$.SC = ($A_{n}$.T // 10) where // is the integer division operator
  • $A_{n}$.SS= ($A_{n}$.T == 9) where == is the equality operator
The A0.T value is accurate. But the other $A_{n}$.T values are not accurate because each is constrained to information from just one digit. We define another more accurate operator $A_{n}$.T2 that has “two-digit accuracy”. $A_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
  • $A_{n}$.T2 = $A_{n}$.T + $A_{n-1}$.SC
$A_{n}$.T2 is more accurate than $A_{n}$.T. The $A_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+CarryOne). Tab. 12 show how the model can implement the T2 operator as a mapping.
$A_{n}$.T $A_{n}$.T2 if $A_{n-1}$.SC==0 $A_{n}$.T2 if $A_{n-1}$.SC==1
0 0 1
1 1 2
9 9 10
10 10 11
17 17 18
18 18 19
Table 12: Calculating $A_{n}$.T2 from $A_{n}$.T and $A_{n-1}$.T
Following this pattern, we define operators $A_{n}$.T3, $A_{n}$.T4 and $A_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
  • $A_{n}$.T3 = $A_{n}$.T + ( $A_{n-1}$.T2 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC2
  • $A_{n}$.T4 = $A_{n}$.T + ( $A_{n-1}$.T3 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC3
  • $A_{n}$.T5 = $A_{n}$.T + ( $A_{n-1}$.T4 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC4
The value A4.T5 is accurate as it integrates SC and cascading SS data all the way back to and including A0.T. The values A1.T2, A2.T3, A3.T4 are also all accurate. If the model knows these values it can calculate answer digits accurately:
  • A1 = A1.T2 % 10
  • A2 = A2.T3 % 10
  • A3 = A3.T4 % 10
  • A4 = A4.T5 % 10
  • A5 = A4.T5 // 10
In this hypothesis, all the answer digits are accurately calculated using the nodes in positions 8 to 11. This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
Experimenation shows the model does not implement this hypothesis. It retains the long staircase SA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
  • Hypothesis 2 is too compact. The model is not optimising for compactness. The long staircase is discovered early in training, and it works for simple questions. Once the overall algorithm gives low loss consistently it stops optimising.
  • Hypothesis 2 accurately predicts all answer digits in step 11 - reflecting the authors (human) values for good code. The model is not motivated to do this. It just needs to accurately predict A5 as 1 or 0 in step 11 and A4 in step 12 - nothing more.
We abandoned this hypothesis.
Appendix J Appendix: Addition Hypothesis 3
The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Some of the experiments and mappings were:
  • Ablation experiments show that the A5 value is accurately calculated in prediction step 11 using 5 attention heads and 5 MLP layers. The pseudo-code accurately calculates A5 while constraining itself to this many steps.
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Table 1). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $A_{n}.ST$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 9).
  • The pseudo-code has 4 instances where compound functions using TriCase and TriAdd to generate tri-state outputs. PCA of the corresponding nodes (P11.L0.H1, P12.L0.H1 and P13.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 9).
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § K for details.
  • The pseudo code includes calculations like A1.ST which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of A1.ST to both nodes.
  • For P10.L0.H1, the attention head PCA could represent either a bi-state or tri-state output (see Figure 10). The MLP layer at P10.L0.MLP could map the attention head output to either a bi-state or tri-state. We cannot see which. The pseudo-code shows a tri-state calculation at P10.L0.MLP, but with small alterations the pseudo-code would work with a bi-state output.
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 10). The pseudo-code shows a bi-state calculation A0.SC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case. The model needs A1.ST2 for A2 accuracy. The model could simply reuse the accurate A1.ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST. That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values”. This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to >99.999% accuracy.
Figure 6: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells. Each cell shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc. The model only uses nodes in nine token positions.
Figure 7: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P8 to P16 (i.e. tokens D’2 to A1). Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 8: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model. At token position P11 the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P11.
Figure 9: For 5-digit addition, for these attention heads, for exactly 1 answer digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA data and TriCase quanta are both tri-state and strongly correlated.
Figure 10: For 5-digit addition, for these attention heads, for exactly 1 digit $A_{n}$each, PCA shows these interpretable results. The dot colours show the TriCase value of each question. The PCA and TriCase data are strongly correlated, but the PCA data has 2 states.
Appendix K Appendix: Addition Interchange Interventions
To test the hypothesis 3 mapping of the mathematical framework (casual abstraction) to the model attention heads, various “interchange interventions” (aka activation patching) experiments were performed on the model, where
  • A particular claim about an attention head has selected for testing.
  • The model predicted answers for sample test questions, and the attention head activations were recorded (stored).
  • The model then predicted answers for more questions, but this time we intervened during the prediction to override the selected attention head activations with the activations from the previous run.
Using this approach we obtained the findings in Tab. 13.
Nodes Claim: Attention head(s) perform … Finding: Attention head(s) perform a function that …
P8.L0.H1 and MLP A2.ST = TriCase( D2, D’2 ) impacting A4 and A5 accuracy Based on D2 and D’2. Triggers on a A2 carry value. Provides carry bit used in A5 and A4 calculation.
P9.L0.H1 and MLP A1.ST = TriCase( D1, D’1 ) impacting A5, A4 and A3 accuracy Based on D1 and D’1. Triggers on a A1 carry value. Provides carry bit used in A5, A4 and A3 calculation.
P10.L0.H1 and MLP A1.ST2 = TriAdd( A1.ST, TriCase(D0, D’0) ) impacting A5, A4, A3 and A2 accuracy Based on D0 and D’0. Triggers on a A0 carry value. Provides carry bit used in A5, A4, A3 and A2 calculation.
P11.L0.H1 and MLP A3.ST4 = TriAdd( TriCase( D3, D’3 ), TriAdd( A2.ST, A1.ST2)) impacting A5 accuracy Based on D3 and D’3. Triggers on a A3 carry value. Provides carry bit used in A5 calculations.
P11.L0.H2 and MLP A4.ST = TriCase(D4, D’4) impacting A5 accuracy A4 Based on D4 and D’4. Triggers on a A4 carry value. Provides carry bit used in A5 calculation.
P12.L0.H0+H2 and MLP A4.SA = (D4 + D’4) % 10 impacting A4 accuracy A4 Sums D4 and D’4. Impacts A4.
P13.L0.H0+H2 and MLP A3.SA = (D3 + D’3) % 10 impacting A3 accuracy Sums D3 and D’3. Impacts A3.
P14.L0.H0+H2 and MLP A2.SA = (D2 + D’2) % 10 impacting A2 accuracy Sums D2 and D’2. Impacts A2.
P14.L0.H1 and MLP (D1 + D’1) / 10 + P10.A1.ST2 info impacting A2 accuracy Calculates P10.A1.ST1 but add P10.A1.ST2 info when A1.ST != A1.ST2. Impacts A2
P15.L0.H0+H2 and MLP A1.SA = (D1 + D’1) % 10 impacting A1 accuracy Sums D1 and D’1. Impacts A1.
P15.L0.H1 and MLP A0.SC = (D0 + D’0) / 10 impacting A1 accuracy Triggers when D0 + D’0 >10. Impacts A1 digit by 1
P16.L0.H0+H2 and MLP A0.SA = (D0 + D’0) % 10 impacting A0 accuracy Sums D0 and D’0. Impacts A0.
Table 13: Interchange Interventions experiments used activation patching to test the claims addition hypothesis 3 made for each attention head in a sample mixed model. Experimental results are consistent with hypothesis 3 for all nodes.
+ve Sub Contains -ve Sub Contains Like
M0 MD N/A N/A S0
M1 MD,MB N1 ND,NB S1
M2 MD,MBx2 N2 ND,NBx2 S2
M3 MD,MBx3 N3 ND,NBx3 S3
M4 MD,MBx4 N4 ND,NBx4 S4
Table 14: We define “positive-answer subtraction” and “negative-answer subtraction” calculation complexity quanta that parallel the addition quanta.
Appendix L Appendix: N-Digit Addition
The addition models perform addition accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
Figure 11: This map shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc.
Figure 12: This map shows the % of questions that fail when we ablate each node in the 6-digit 2-layer 3-head addition model. The model only uses nodes in token positions P11 to P20. Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not useful.
Some notes about the models:
  • The models selected different attention heads in the early positions to use to do the same logical calculations.
  • Some models use 2 attention heads per digit to do the SA calculation, whereas some models only uses one (and so are more compact).
  • The PCA trigrams have difference appearances in different models (but the same interpretable clusters). Refer Figures 9
  • Per answer digit, some models use the SC calculation, whereas some models optimize it out and rely solely on the $ST$ value (and so are more compact).
Appendix M Appendix: Mixed Model Initialization
We experimented with three approaches to re-using the trained addition model in the “mixed” (addition and subtraction) model:
  • Initialize Only: Initialize the untrained mixed model with the addition model weights before training begins.
  • Freeze Attention: As per “Initialize Only”, but also every 100 training steps recopy the attention head weights from the addition model into the partially-trained mixed model.
  • Freeze All: As per “Initialize Only”, but also every 100 training steps recopy the entire addition model (attention heads and MLP layers) into the partially-trained mixed model.
Our intuition was that “Initialize Only” would give the mixed model the most freedom to learn new algorithms, but that the “Freeze Attention” and “Freeze All” approaches would make the resulting trained mixed model easier to interpret (as we could reuse our addition model insights).
After experimentation we found that the “Initialize Only” approach was the only one that quickly trained to be able to do both addition and subtraction accurately. We concluded that the other two methods constrain the model’s ability to learn new algorithms too much.
We also experimented with “where” in the model we inserted the addition (6-digit, 2-layer, 3-head) model into the slightly larger (6-digit, 3-layer, 4-head) mixed model. That is, do we initialize the first 2 layers or the last 2 layers of the mixed model? Also do we initialize the first 3 attention heads or the last 3 attention heads of the mixed model? Our intuition was that initializing the first layers and heads would be more likely to cause the model to re-use the addition circuits adding interpretability, so we used this approach.
Appendix N Appendix: N-Digit Subtraction
The mixed models perform addition and subtraction accurately. Visualizations that provided insights into the behavior of the model, aiding our interpretation of the algorithm, are below:
Figure 13: This map of a sample 6-digit mixed model shows the 98 nodes used to predict answers to addition (S), positive-answer subtraction (M) and/or negative-answer subtraction (N) questions. Before training the mixed model, 48 nodes were initalized pre-training with a smaller addition model’s weights. These are have a red border. During mixed model training, 39 of 48 of the initalized monosemantic nodes were generalized (become poly-semantic) and now help predict two or three question classes.
Figure 14: This map shows the simplest complexity quanta S0, S1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model when doing addition questions.
Figure 15: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers.
Some notes about the mixed models:
  • All the notes about the addition model (above) also apply to the mixed model.
  • The model contains a new subtask that stands out: The algorithm relies on calculations done at token position P0, when the model has only seen one question token! What information can the model gather from just the first token? Intuitively, if the first token is a “8” or “9” then the first answer token is more likely to be a “+” (and not a “-”). The model uses this heuristic even though this probabilistic information is sometimes incorrect and so will work against the model achieving very low loss.